### v2

# Region-Wise Attentive Multi-View Representation Learning For Urban Region Embedding

Abstract.
Urban region embedding is an important and yet highly challenging issue due to the complexity and constantly changing nature of urban data. To address the challenges, we propose a Region-Wise Multi-View Representation Learning (ROMER) to capture multi-view dependencies and learn expressive representations of urban regions without the constraints of rigid neighborhood region conditions. Our model focuses on learning urban region representation from multi-source urban data. First, we capture the multi-view correlations from mobility flow patterns, POI semantics, and check-in dynamics. Then, we adopt global graph attention networks to learn the similarity of any two vertices in graphs. To comprehensively consider and share features of multiple views, a two-stage fusion module is further proposed to learn weights with external attention to fuse multi-view embeddings. Extensive experiments for two downstream tasks on real-world datasets demonstrate that our model outperforms state-of-the-art methods by up to 17% improvement.
Graph Attention, Multi-Feature Fusion, Urban Region Embedding, Graph Neural Network
journalyear: 2023copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Information and Knowledge Management; October 21–25, 2023; Birmingham, United Kingdombooktitle: Proceedings of the 32nd ACM International Conference on Information and Knowledge Management (CIKM ’23), October 21–25, 2023, Birmingham, United Kingdomprice: 15.00doi: 10.1145/3583780.3615194isbn: 979-8-4007-0124-5/23/10submissionid: shp2626ccs: Computing methodologies Neural networksccs: Information systems Clusteringccs: Information systems Data analytics
1. Introduction
Urban region embedding is a classical embedding problem whose purpose is to learn quantitative representations of regions from multi-sourced data. The problem has been found useful in many real-world applications such as socio-demographic feature prediction (Jean et al., 2018; Wang et al., 2020), crime prediction (Wang and Li, 2017; Zhang et al., 2020; Wu et al., 2022), economic growth prediction(Hui et al., 2020) and land usage classification (Yao et al., 2018; Zhang et al., 2020; Wu et al., 2022; Luo et al., 2022b).
Urban region embedding is challenging due to the highly complex nature of urban data, deep learning approaches have attracted much interest in recent years. In particular, GNNs have shown great potential for learning low-dimensional embeddings of graph-structured data (Cui et al., 2017; Abu-El-Haija et al., 2018; Kipf and Welling, 2016; Gilmer et al., 2017; Wang et al., 2021b, a). Several existing studies that have attempted to integrate POI data and human mobility to characterize regions and obtain attractive results(Zhang et al., 2017; Wang and Li, 2017; Fu et al., 2019; Yao et al., 2018; Zhang et al., 2020; Chang et al., 2020; Liu et al., 2019; Wu et al., 2022; Luo et al., 2022a). For example, the methods (Zhang et al., 2020; Wu et al., 2022) learns graph embeddings by combining multi-graph. The method(Zhang et al., 2020) employs an attention mechanism and simple crosstalk operations to aggregate information, while MFGN(Wu et al., 2022) deeply resolves the relevance of regions in fine-grained human mobility patterns. However, existing studies have predominantly focused on capturing correlations between neighboring regions, overlooking the significant influence of distant regions in complex urban systems. Attention mechanism such as GAT(Veličković et al., 2018) assigns aggregation weights based solely on the influence of neighboring nodes, and self-attention(Vaswani et al., 2017) solely focuses on the regions themselves, both approaches fail to consider the potential correlations between regions and suffer from computational complexity(Guo et al., 2021). As a result, these methods produce sub-optimal embeddings and limit the ability of the model to capture the underlying urban dynamics and features.
To tackle these challenges, we propose a Region-Wise Multi-Graph Representation Learning (ROMER) for effective urban region representation with multi-view data. Our ROMER model adopts a heterogeneous graph neural framework with respect to human movement patterns, POI semantics, and check-in dynamics. In addition, a multi-graph aggregation module is designed to capture region-wise dependencies and non-linear correlations among regions. Finally, we design an efficient and cost-effective attentive fusion module that learns adaptive weights for information sharing across diverse views with external attention and gating mechanisms to fuse multi-view in an efficient and deeply collaborative manner.
We have extensively evaluated our approach through a series of experiments using real-world data. The results demonstrate the superiority of our method compared to state-of-the-art baselines, achieving improvements of up to 17%. In addition, our method exhibits significant computational efficiency.
2. Preliminaries
In this section, we first give some notations and define the urban region embedding problem. We partition a city into $N$ regions $V=\{v_{1},v_{2},\cdots,v_{N}\}$, where $v_{i}$ denotes the $i$-th region. For a trip $p=(v_{o},v_{d})$, $v_{o}$ denotes the origin region and $v_{d}$ denotes the destination region where $1\leq o,d,\leq N$. Given the set of regions $V$, we further define the following.
Definition 1. (Human mobility feature). The human mobility feature is defined as a trip set $\mathcal{P}=\{p_{i}|p_{i}.{v_{o}},p_{i}.{v_{d}}\in V\}$ that occur in urban areas. $i=\{1,2,\cdots,M\}$, where $M$ is the number of trips.
Definition 2. (Semantic feature). The semantic feature describes the functional similarity among regions. Similar regions may necessarily be close in space. In this work, the semantic feature of the region is characterized by POIs in the located region. Given a region $v_{i}$, its semantic feature is represented as:
(1) $\mathcal{S}=\left\{s_{i}\mid s_{i}\in\mathbb{R}\right\},i=1,2,\cdots,n$
Definition 3. (Dynamic feature). The dynamic feature describes the activities of POIs in regions, which integrates human activities and POIs information. Given a region $v_{i}$, its dynamic feature is represented as:
(2) $\mathcal{G}=\left\{g_{i}\mid g_{i}\in\mathbb{R}\right\},i=1,2,\cdots,m$
Urban region embedding problem. We denote the three families of features for a region $v_{i}$ as a vector $e^{i}\in\mathbb{R}^{r}$, where $r$ is the number of features. Then, our final goal is to learn a mapping function $F$,
(3) $\mathcal{E}=F(\mathcal{E^{V}})$
where $\mathcal{E^{V}}\in\mathbb{R}^{M\times r}$ are three features of all regions in $V$.
3. METHODOLOGY
The overall architecture of ROMER proposed in this paper is shown in Figure. 1, which consists of three components: the region-wise graph learning module, the multi-graph aggregation module, and the attentive fusion module.
3.1. Region Wise Graph Learning Module
In this section, we elucidate the utilization of various types of region dependencies for encoding multi-graph.
3.1.1. Mobility-based Region Graph
The movement of people within urban spaces across regions can be understood by examining their interactions. When people travel between different origins ($O$) and destinations ($D$), we can observe similarities in their patterns if they have the same O/D region. In simpler terms, by analyzing the similarity of O/D patterns, we can identify important potential features related to human mobility. Given a set of human mobility $M$, we can use
(4) $s_{v_{j}}^{v_{i}}=\left|\left(v_{i},v_{j}\right)\in M\right|$
to calculate the similarity value between region $v_{o}$ and region $v_{d}$, where |.||.|| . | counts the size of the trip. Then we employ distributions $p_{o}\left(v\mid v_{i}\right)$ and $p_{d}\left(v\mid v_{i}\right)$ to describe the origin and destination contexts of a region $v_{i}$ as follows:
(5) $p_{o}\left(v\mid v_{i}\right)=\frac{s_{v_{i}}^{v}}{\sum_{v}s_{v_{i}}^{v}},\quad p_{d}\left(v\mid v_{i}\right)=\frac{s_{v}^{v_{i}}}{\sum_{v}s_{v}^{v_{i}}}.$
The two types of dependencies were defined by us based on the source and destination context of each region, as follows,
(6) $\mathcal{D}_{O}^{ij}=\operatorname{sim}\left(p_{o}\left(v\mid v_{i}\right),p_{o}\left(v\mid v_{j}\right)\right),$
(7) $\mathcal{D}_{D}^{ij}=\operatorname{sim}\left(p_{d}\left(v\mid v_{i}\right),p_{d}\left(v\mid v_{j}\right)\right),$
where $\mathcal{D}_{O}^{ij}$ denotes the dependencies between two origins, $\mathcal{D}_{D}^{ij}$ represents the dependencies between two destinations, $\operatorname{sim}(\cdot)$ denotes the cosine similarity. Based on Equations (6) and (7), we construct region-wise graphs $\mathcal{G}_{O}$ and $\mathcal{G}_{D}$.
3.1.2. Semantic Region Graph
In urban environments, the description of regional semantics relies on the utilization of Point of Interest (PoI) information. The PoI attributes encapsulate the semantic features associated with specific regions. To incorporate the PoI context into region embeddings, we leverage semantic dependencies to effectively capture and integrate region functionality information into the representation space. The calculation for this process can be described as follows:
(8) $\vspace{-0.5em}\mathcal{D}_{S}^{ij}=\operatorname{sim}\left(\vec{s}_{i},\vec{s}_{j}\right).$
where $\mathcal{D}_{S}^{ij}$ is the semantic dependency between region $v_{i}$ and $v_{j}$. We thus obtain the semantic region graph $\mathcal{G}_{S}$.
3.1.3. Dynamic Region Graph
In contrast to PoI attributes, which solely provide information regarding the quantity of PoIs, check-in data takes into account human activity and reflects the popularity of each PoI category. When characterizing regions with check-in attributes, we employ a dynamic dependency measure to determine the significance of each check-in type within a given region. The calculation of this measure can be described as follows:
(9) $\vspace{-0.5em}\mathcal{D}_{G}^{ij}=\operatorname{sim}\left(\vec{a}_{i},\vec{a}_{j}\right)$
where $\mathcal{D}_{G}^{ij}$ is the dynamic dependency between region $v_{i}$ and $v_{j}$. Now we construct dynamic region-wise graph $\mathcal{G}_{G}$.
Figure 1. The architecture of ROMER consists of three key components: A) the region-wise graph learning module, B) the multi-graph aggregation module, and C) the attentive fusion module.
3.2. Multi-Graph Aggregation Module
It is observed that not only adjacent regions are relevant, but also many distant regions are correlated. However, existing GAT-based methods (Zhang et al., 2020; Luo et al., 2022a) only consider the influence of neighboring nodes. Inspired by (Huang et al., 2020), we utilize an improved GAT mechanism to extract any two relevant regions in a city to assign learning weights. Given the vertex feature $\mathbf{h}=\left\{\vec{h}_{1},\vec{h}_{2},\ldots,\vec{h}_{n}\right\},\vec{h}_{i}\in\mathbb{R}^{F}$, where $F$ is the input dimension, the MGA layer works as follows:
(10) $\begin{split}&A_{ij}=cosine(h_{i},h_{j})\cdot w_{ij}=\frac{h_{i}(h_{j})^{T}\cdot w_{ij}}{\lVert h_{i}\rVert\lVert h_{j}\rVert},\\ &\hat{A_{ij}}=\sigma(A_{ij}),\end{split}$
where $\sigma$ denotes the softmax function, $w_{ij}$ is the weight matrix. $A_{ij}$ is the similarity between $h_{i}$ and $h_{j}$. The softmax function is used to normalize the coefficients. $\lVert h_{i}\rVert$ denotes the norm of vector $h_{i}^{l}$ and $\cdot$ is the dot product of vectors.
Next, $\hat{A_{ij}}$ is utilized to aggregate information from all other features in the network to each feature, by
(11) $\hat{h_{i}}=\sum_{j\in\widetilde{N}_{i}}\hat{A_{ij}}h_{j}w_{ij},$
where $\widetilde{N}_{i}$ is all nodes in the graph except node $h_{i}$. $\hat{h_{i}}$ is the information aggregation from the global features to the feature $h_{i}$.
In our model, $\mathcal{G}_{O}$, $\mathcal{G}_{D}$, $\mathcal{G}_{S}$ and $\mathcal{G}_{G}$ are fed into the MAG block, then we obtain the corresponding representation results as $\mathcal{E}_{O}$, $\mathcal{E}_{D}$, $\mathcal{E}_{S}$, and $\mathcal{E}_{G}$.
Table 1. Performance comparison of different approaches for check-in prediction and land usage classification.
Models Check-in Prediction Land Usage Classification
MAE RMSE $R^{2}$ NMI ARI
LINE 564.59 853.82 0.08 0.17 0.01
node2vec 372.83 609.47 0.44 0.58 0.35
HDGE 399.28 536.27 0.57 0.59 0.29
ZE-Mob 360.71 592.92 0.47 0.61 0.39
MV-PN 476.14 784.25 0.08 0.38 0.16
MVGRE 297.72 495.27 0.63 0.78 0.59
MGFN 280.91 436.58 0.72 0.76 0.58
ROMER(ours) 252.14 413.96 0.74 0.81 0.68
3.3. Attentive Fusion Module
In this section, we design an attentive fusion mechanism to efficiently interchange information among multiple views. The self-attention mechanism is widely used in existing fusion methods(Vaswani et al., 2017), which is at the price of high computation. Inspired by (Guo et al., 2021), We adopt external attention to allow information to propagate across multiple views. Given the representations of $M$ views $\{\mathcal{E}_{1},\mathcal{E}_{2},\cdots,\mathcal{E}_{M}\}$, for each $\mathcal{E}_{i}$, we then propagate information among all views as follows:
(12) $\left[A_{i}\right]_{i=1}^{M}=\operatorname{Norm}\left(\left[{\mathcal{E}_{i}M_{k}^{T}}\right]_{i=1}^{M}\right),\hat{\mathcal{E}}_{i}=\sum_{i=1}^{M}A_{i}M_{v}$
where $M_{k}\in\mathbb{R}^{S\times d}$ and $M_{v}\in\mathbb{R}^{S\times d}$ are learnable parameters independent of the representations, which act as the key and the value memory of the whole training dataset. And $d$ and $S$ are hyper-parameters. ^$\hat{\mathcal{E}}_{i}$ is considered as the relevant global information for $i$-th view. Meanwhile, the normalization method we follow (Guo et al., 2021). The embedding results ^$\hat{\mathcal{E}}_{O}$, ^$\hat{\mathcal{E}}_{D}$, ^$\hat{\mathcal{E}}_{S}$, and ^$\hat{\mathcal{E}}_{G}$ are generated from the above modules, linking the global information to subsequent fusions in the model.
To integrate global and local region representation. We follow the fusion mechanism in (Zhang et al., 2020). The fusion layer operates as follows:
(13) $\mathcal{E}_{i}^{\prime}=\alpha\hat{\mathcal{E}}_{i}+(1-\alpha)\mathcal{E}_{i},\quad 0\leq\alpha\leq 1$
(14) $\mathcal{E_{F}}=\sum_{i}^{M}w_{i}\mathcal{E}_{i},w_{i}=\sigma\left(\mathcal{E}_{i}W_{f}+b_{f}\right)$
where $\mathcal{E}_{i}^{\prime}$ is the representation for $i$-th view with global information, and $\alpha$ is the weight of global information, $wi$ is the weight of $i$-th view, which is learned by a single layer MLP network with the $i$-th embeddings as input.
In a bid to enable the learning of the multi-view fusion layer, we engage $\mathcal{E}$ in the learning objective of each view. Formally, we update the representation of each view as:
(15) ~$\tilde{\mathcal{E}}_{i}=\left(\mathcal{E}_{i}^{\prime}+\mathcal{E_{F}}\right)/2.$
By incorporating the outputs of the base model into the proposed joint learning module, we derive region embeddings ~$\tilde{\mathcal{E}}_{O},\tilde{\mathcal{E}}_{D},\tilde{\mathcal{E}}_{S}$, and ~$\tilde{\mathcal{E}}_{G}$, on which we work out the various learning goals.
Figure 2. Ablation studies for two tasks on NYC dataset. (a) MAE and RMSE in Check-in Prediction. (b) ARI and NMI in Land Usage Classification.
3.4. Prediction Objectives
In order to efficiently train our model, according to (Zhang et al., 2020), given the source region $v_{i}$, we model the distribution of the target region $v_{j}$ as follows:
(16) $\hat{p}_{O}\left(v_{j}\mid v_{i}\right)=\frac{\exp\left(e_{O}^{iT}e_{D}^{j}\right)}{\sum_{j}\exp\left(e_{O}^{i}e_{D}^{j}\right)}.$
Similarly, we model $\hat{p}_{D}\left(v_{j}\mid v_{i}\right)$ similarly for the distribution of source region $r_{i}$ for a given destination region $r_{j}$. Then, $\mathcal{L}_{HMF}$ is constructed by maximizing the probability of O/D occurrence. Hence, the $\mathcal{L}_{HMF}$ between region $r_{i}$ and region $r_{j}$ can be computed as:
(17) $\mathcal{L}_{HMF}=\sum_{\left(v_{i},v_{j}\right)\in\mathcal{M}}-\log\hat{p}_{o}\left(v_{j}\mid v_{i}\right)-\log\hat{p}_{d}\left(v_{i}\mid v_{j}\right).$
For the learned region embeddings to preserve the region similarity across region attributes, we designed tasks to reconstruct region correlations based on the corresponding embeddings. Taking the Check-in property as an example, the learning objective is based on $\mathcal{C}_{G}$ and ~$\tilde{\mathcal{E}}_{G}=\left\{e_{G}^{i}\right\}_{i=1}^{n}$ defined as follows.
(18) $\mathcal{L}_{VF}=\sum_{i,j}\left(\mathcal{C}_{G}^{ij}-e_{G}^{i{}^{T}}e_{G}^{j}\right)^{2}.$
Similarly, we define the learning objective $\mathcal{L}_{SF}$ of check-in attributes. In this way, the final learning objective is:
(19) $\mathcal{L}=\mathcal{L}_{HMF}+\mathcal{L}_{VF}+\mathcal{L}_{SF}.$
4. EXPERIMENTS
In this section, extensive experiments are conducted to verify the superiority of the proposed model.
4.1. Datasets and Baselines
Experiments are conducted on several real-world datasets of New York City from NYC open data website [^1] https://opendata.cityofnewyork.us/. We follow (Zhang et al., 2020), taxi trip data was analyzed as the source of human movement data, and the borough of Manhattan was divided into 180 zones to serve as the study area. As shown in Figure.  3(a), the borough of Manhattan was divided into 12 regions based on land use according to community board (Berg, 2007).
This paper compares the ROMER model with the following region representation methods, including LINE(Tang et al., 2015), node2vec(Grover and Leskovec, 2016), HDGE(Wang and Li, 2017), ZE-Mob(Yao et al., 2018), MV-PN(Fu et al., 2019), MVGRE(Zhang et al., 2020), MGFN(Wu et al., 2022) etc.
(a) Districts
(b) HDGE
(c) ZE-Mob
(d) MGFN
(e) MVGRE
(f) Ours
Figure 3. Districts in Manhattan and region clusters.
4.2. Experimental Settings
For land usage classification, we clustered the region embeddings using K-means with Normalized Mutual Information (NMI) and Adjusted Rand Index (ARI) with settings (Yao et al., 2018).In the case of check-in prediction, we utilize the Lasso regression (Tibshirani, 1996) with metrics of Mean Absolute Error (MAE), Root Mean Square Error (RMSE), and coefficient of determination ($R^{2}$).
4.3. Experimental Results
4.3.1. Main Results
Table. 1 shows the results of the check-in prediction task and the land usage classification task. We draw the following conclusions: (1) Our method (ROMER) outperforms all baseline tasks, in particular achieving over 10% improvement in MAE in the check-in prediction task and over 17% improvement in ARI in the land usage classification task. (2) Traditional graph embedding methods (LINE, node2vec) perform poorly because of the local sampling approach, which may not fully express the relationships between nodes. (3) While HDGE, ZE-Mob, and MV-PN utilize multi-scale graph structures and embedding methods to capture the multi-level features and complex relationships within the urban system, they lack an attention mechanism that accounts for the varying importance of nodes. (4) Both MVGRE and MGFN employ multi-view fusion methods and attention mechanisms. However, the long-range dependence of the regions mined by these models is poor.
4.3.2. Ablation Study
To verify the effect of key components on the proposed model, this paper conducts an ablation study in land usage classification and check-in prediction tasks respectively. The variants of ROMER are named as follows:
  • ROMER-G: It is ROMER without a multi-graph aggregation Module (MGA). MGA is replaced with the GAT(Veličković et al., 2018).
  • ROMER-A: It is ROMER without Attentive Fusion which is replaced with the self-attention(Vaswani et al., 2017).
  • ROMER-D: It is ROMER without Attentive graph fusion. The extracted spatial features are concatenated directly.
Figure. 2 displays the experimental results of ROMER and its variants in check-in prediction and land use classification tasks. The MGA module (ROMER-G) exhibits the most substantial impact on performance. Without the MGA component, the MAE for the prediction task increases significantly from 252.14 to 351.52, and the RMSE increases from 413.96 to 559.98. The attentive fusion module (ROMER-A) demonstrates the second-largest impact, affirming the effectiveness of our constructed module in enhancing single-view performance.
4.3.3. Visualized Analysis
To visually evaluate the clustering results, we plotted the clustering results of five methods in Figure. 3, where the same color marks the regions in the same cluster. We observe that the clustering results based on our method are optimal in terms of consistency with the real boundaries of ground conditions. These results suggest that the regional embeddings learned by our model can represent regional functions effectively.
5. Conclusion
In this paper, we try to solve the urban region embedding problem with attentive multi-view neural networks. Specifically, we designed a graph aggregation module to capture region-wise dependencies within the urban network. To comprehensively share information across multiple views, we designed an attentive fusion module and fuse view embeddings with external attention and gating mechanisms. Extensive experimental results on real-world data demonstrated the effectiveness of our proposed ROMER. In the future, we will apply the proposed framework to additional graph-based applications.
Acknowledgments
This work was supported by the China Postdoctoral Science Foundation under Grant No.2022M711088.